# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v3

@@ 3.2 Convergence time
The proof of this lemma is provided in Appendix B.3. It provides a bound on $\rho(PP^{*})$ for any transition matrix that satisfies Equation (7). It is in particular applicable to the Markov chain induced by log-linear learning since the transition matrix specified in Equation (4) satisfies Equation (7).
Combination: Combining the two parts of the proof we deduce that:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\mathbb{E}_{a\sim\mu}[\Phi(a)]-2\|\mu^{t}-\mu\|_{TV}\max_{a\in\mathcal{A}^{N}}\Phi(a)$
$\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\frac{\epsilon}{2}-\frac{2\epsilon}{4},$
for $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)$. This concludes the proof of the theorem.
In the following, we additionally assume that the potential game is symmetric, i.e., players are interchangeable. [^6] A definition of a symmetric game is given in Appendix B.4, see Definition B.6. In these games, a player’s utility depends solely on the number of players selecting each action and not on their identity. Technically speaking, in this case, log-linear learning induces a Markov chain on a modified state space of cardinality $\mathcal{O}(N^{A})$, see Equation (22), rather than on the state space $A^{N}$. Exploiting the smaller state space, we derive a bound on the convergence time of log-linear learning that depends polynomially not only on $1/\epsilon$ but also on $N$.
Corollary 3.5 (Sketch).
Consider a symmetric potential game. For any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to modified log-linear learning (Shah and Shin, 2010) with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log(\frac{N^{A}}{\epsilon}))$. Then, $\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for $t=\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$.
We provide a full statement of the corollary and its proof in Appendix B.4. For symmetric potential games with a $\lambda$-Lipschitz-continuous potential function, Shah and Shin (2010) proves a convergence time of $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$. In comparison, relaxing the Lipschitzness assumption comes at a small cost of polynomial dependence on $N$ rather than a linear one. However, our result greatly improves the dependence on $\epsilon$ to a polynomial one.
4 Robustness of log-linear learning
In this section, we show that our convergence time guarantee is robust to settings where players have less feedback information and to settings where log-linear learning is subject to perturbations such as noisy utility observations.
4.1 Reduced feedback
Log-linear learning requires players to access their utilities for all possible actions given the other players’ actions. Having such full-information feedback when action sets are large can be demanding. Binary log-linear learning (Arslan et al., 2007; Marden et al., 2007) alleviates this limitation by requiring two-point feedback, reducing the feedback needed by a factor $A$ per round. Now, we briefly review the binary-log-linear learning rule. Then, we derive the first finite-time convergence bound of binary log-linear learning to an $\epsilon$-efficient Nash equilibrium, showing that the deterioration in convergence time is a constant.
$p_{i}^{t}(a)=\begin{cases}\frac{e^{\beta U_{i}(a,a_{-i}^{t-1})}}{e^{\beta U_{i}(a_{i}^{t-1},a_{-i}^{t-1})}+e^{\beta U_{i}(\tilde{a}_{i},a_{-i}^{t-1})}}&\text{ for }a\in\{a_{i}^{t-1},\tilde{a}_{i}\}\\ 0&\text{ otherwise}\end{cases}$
Here, player $i$ can either repeat her action $a_{i}^{t-1}$ or play one other randomly sampled action $\tilde{a}_{i}$ rather than any action $a_{i}\in\mathcal{A}$ as in log-linear learning.
Theorem 4.1.
Consider a potential game with potential function $\Phi:\mathcal{A}^{N}\rightarrow[0,1]$ and $A\geq 4$. For any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to binary log-linear learning with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{\mathcal{A}^{N}}{\epsilon}).$ Then, it holds that $\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for
$t\geq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)\approx\tilde{\mathcal{O}}\left(N^{2}A^{5}\left(\frac{A^{N}}{\epsilon}\right)^{\frac{N}{\max\{\epsilon,\Delta\}}}\right).$
Theorem 4.1 shows the first finite-time convergence guarantee for binary log-linear learning, we provide its detailed proof in Appendix C. It is remarkable that with significantly less feedback per round, binary log-linear achieves the same convergence speed as log-linear learning up to a factor of $8$ (Theorem 3.1). This raises the question of whether two-point feedback is sufficient for learning $\epsilon$-efficient Nash equilibria or whether the convergence time bounds we proved for log-linear learning are loose.
4.2 Perturbed log-linear learning
Classical log-linear learning relies on two limiting assumptions: 1) Players have access to their exact utilities. However, in real-world applications, the presence of noise is typical as uncertainties and hidden factors generate inexact measurements. 2) Players are rational. However, empirical evidence suggests that players have limited rationality and therefore may occasionally deviate from the log-linear learning rule in practical scenarios. Our next result generalizes Theorem 3.1 to the case where the log-linear learning rule is subject to small perturbations. As we will show, this generalization can address the two limitations above.
We first derive a general statement for Markov chains which shows that the induced stationary distribution is Lipschitz-continuous as a function of the transition matrix. We then use this lemma to prove our main result on the convergence time of perturbed log-linear learning to an $\epsilon$-efficient NE.
Lemma 4.2 (Lipschitzness).
Consider two irreducible and aperiodic transition matrices $P_{1},P_{2}\in\mathbb{R}^{A^{N}\times A^{N}}$. Let $\mu_{1}$ and $\mu_{2}$ be the stationary distributions of the Markov chains induced by $P_{1}$ and $P_{2}$, respectively. Then, the following holds:
$\|\mu_{1}-\mu_{2}\|_{2}\leq\min\{L(P_{1}),L(P_{2})\}\|P_{1}-P_{2}\|_{2},$
We provide a proof in Appendix D.4. Compared to the result of (Zhang et al., 2023), we considerably improve the Lipschitz constant by using the mixing time bound based on log-Sobolev inequalities (Lemmas 2.2 and 3.4). In particular, (Zhang et al., 2023, Lemma 24) entails a Lipschitz constant $L=\tilde{\mathcal{O}}((e/p_{\min})^{N})$ while Lemma 4.2 implies that $L=\tilde{\mathcal{O}}(1/(\mu_{\min}p_{\min}^{3}))$. [^7] This can be seen by injecting the log-Sobolev bound of lemma 3.4 into our Lemma above. Now, we state the main result of this section.
Theorem 4.3.
Consider a potential game with a potential function $\Phi:\mathcal{A}^{N}\rightarrow[0,1]$ and $A\geq 4$. Furthermore, consider a learning rule with transition matrix $P$ and assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that for all $a,\tilde{a}\in\mathcal{A}$ it holds that:
$\frac{1}{N}p_{\min}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}\leq P_{a,\tilde{a}}\leq\min\{1,\frac{1}{N}p_{\max}\}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}.$ (8)
For any $\epsilon\in(0,1)$ and any initial distribution $\mu^{0}$, assume all players adhere to this learning rule with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$. Then,
for
$t\geq\frac{25N^{3/2}e^{N}}{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}\log\left(\frac{4A^{N}}{\epsilon^{2}}\log\frac{e^{N}}{p_{\min}^{N}\sqrt{2\pi N}}\right),$
where $P_{\ell}$ is the transition matrix of log-linear learning and $L$ is a Lipschitz constant of order $\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
Theorem 4.3 shows that small perturbations of the log-linear learning rule do not compromise the convergence to an $\epsilon$-efficient NE. In particular, if the players follow a learning rule $P$ with $\|P-P_{\ell}\|_{2}=\mathcal{O}(\epsilon/(L\sqrt{A^{N}}))$ then they converge to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$. On the other hand, due to the unavailability of the stationary distribution of the perturbed learning rule, we suffer an extra factor of $(N/p_{\min})^{N}/N!$ in the convergence time guarantee compared to log-linear learning.
4.2.1 Corrupted utilities with additive noise
In the following, we assume that players observe noise-corrupted utilities $(\hat{U}_{i})_{i\in[N]}$ which satisfy:
where $\xi_{i}(a_{i},a_{-i})\in[-\xi,\xi]$ is a bounded noise term. Alternatively, the noise could be assumed to be centered i.i.d. random variables with bounded variance (Leslie and Marden, 2011). Using Theorem 4.3, we hereafter prove that log-linear learning is robust to noisy feedback.
Corollary 4.4.
Consider the setting of Theorem 4.3 with noise-corrupted utilities as in Equation (9). If all players adhere to log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$ and $\xi\leq 1/(2\beta)$, then
for $t=\mathcal{O}\left(N^{3/2}A^{3}e^{N+\beta(1+2\xi)(N+3)}\log\frac{1}{\epsilon^{2}}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
The proof is provided in Appendix D.2. Corollary 4.4 shows that log-linear learning with corrupted utilities converges to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$ if the corruption magnitude $\xi$ is sufficiently small. Our finite-time convergence result extends previous works on robust learning which provide asymptotic guarantees (Leslie and Marden, 2011; Lim and Shamma, 2013; Bravo and Mertikopoulos, 2017). The key to this result lies in showing that the transition matrix of the Markov chain induced by corrupted utilities is close to its corruption-free counterpart.
4.2.2 Log-linear learning mixed with uniform exploration
In the following, we assume players occasionally explore actions randomly. A modification of log-linear learning based on the fixed-share algorithm (Herbster and Warmuth, 1998) can reflect such a random behavior. In the so-called fixed-share log-linear learning, a player $i$ is randomly chosen and allowed to alter her action. Player $i$ samples her new action from the following distribution:
The exploration parameter $\xi\in(0,1)$ determines how likely a player is to act randomly, where a value of $\xi=1$ corresponds to a uniform action sampling while $\xi=0$ corresponds to standard log-linear learning. For simplicity, we focus on the full-information case but fixed-share log-linear learning can easily be adapted to the binary setting. Note that this modification resembles the $\epsilon$-Hedge strategy (Heliou et al., 2017) in the expert advice literature, and under binary feedback, this modification resembles the Epx3.P strategy (Auer et al., 2002; Bubeck et al., 2012) in the multi-armed bandit literature. Here, the fixed share $\xi/A$ ensures a lower bound on the exploration.
Without explicit knowledge of the stationary distribution of this learning rule, we can apply Theorem 4.3 to deduce the following result.
Corollary 4.5.
Consider the setting of Theorem 4.3, where all players adhere to fixed-share log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$. Then, for any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$ we have:
for $t=\mathcal{O}\left(N^{3/2}A^{N+3}e^{\beta(N+3)}/(1-\xi)^{N+3}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
We provide a proof in Appendix D.3. Corollary 4.5 guarantees the convergence of fixed-share log-linear learning to an $\epsilon$-efficient Nash equilibrium in time polynomial in $1/\epsilon$ if the exploration parameter $\xi$ is sufficiently small. The key to this result is to show that the transition matrix of fixed-share log-linear learning is close to the transition matrix of the unperturbed learning rule in terms of the $\ell_{2}$ distance.
5 Conclusion
We provided the first finite-time convergence guarantees to an $\epsilon$-efficient NE for potential games using a novel mixing-time bound based on a log-Sobolev constant. In particular, we guarantee a polynomial dependence on $1/\epsilon$ using a problem-dependent analysis. Furthermore, under the additional assumption that the game is symmetric, we showed that the exponential dependence on the number of players $N$ present in our bound on the convergence time can be avoided. To deal with reduced feedback, i.e., two-point feedback on the utility, we considered binary log-linear learning and showed that it enjoys the same convergence time as log-linear learning up to numerical constants. Finally, we proved that the convergence time of log-linear is not hindered by corruptions of the observed utilities by bounded noise or by small perturbations in the learning rule. The relevance of this result is twofold: First, our analysis does not rely on characterizing the stationary distribution of this perturbed Markov chain; Second, the presence of noise is ubiquitous in real-world applications and it is therefore crucial that the implemented learning rule is robust to such corruptions.
In future work, we are interested in providing lower bounds on the convergence time of log-linear learning to an $\epsilon$-efficient NE. Such bounds would allow us to assess the tightness of our results.
Appendix A Markov chains and mixing times
In the following, we define the log-Sobolev constant and the spectral gap of a Markov chain. We then discuss the implications of Lemma 2.2 and provide a proof of the lemma.
Definition A.1 ((Diaconis and Saloff-Coste, 1996)).
Consider a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ over state space $\mathcal{A}^{N}$ with transition matrix $P$ and stationary distribution $\mu$. The log-Sobolev constant $\rho(P)$ is defined as:
$\rho(P):=\inf_{\mathcal{L}_{\pi}(f^{2})\neq 0}\frac{\mathcal{E}_{P}(f,f)}{\mathcal{L}_{\pi}(f^{2})}.$ (10)
The spectral gap $\lambda(P)$ is defined as:
$\lambda(P):=\inf_{\text{Var}_{\pi}(f)\neq 0}\frac{\mathcal{E}_{P}(f,f)}{\text{Var}_{\pi}(f)}.$ (11)
For any $f:\mathcal{A}^{N}\rightarrow\mathbb{R}$, the Dirichlet form $\mathcal{E}_{P}(f,f)$ is defined by:
the entropy-like quantity $\mathcal{L}(f^{2})$ is defined by:
$\mathcal{L}(f^{2})=\sum_{a\in A^{N}}f(a)^{2}\log\frac{f(a)^{2}}{\|f\|_{2}^{2}}\mu(a),$
and the variance $\text{Var}_{\pi}(f)$ is defined by:
$\text{Var}_{\pi}(f)=\sum_{a,\tilde{a}\in A^{N}}(f(a)-f(\tilde{a}))^{2}\mu(a)\mu(\tilde{a}).$
As mentioned in Section 2.2, using the log-Sobolev constant $\rho(PP^{*})$ can often significantly improve classical mixing time bounds based on the spectral gap $\lambda(PP^{*})$. Such classical bounds are of the form (Montenegro et al., 2006):
where $C$ is some constant. Diaconis and Saloff-Coste (1996) in Lemma $3.1$ show that the log-Sobolev constant $\rho(PP^{*})$ is upper-bounded by the spectral gap $\lambda(PP^{*})$ as follows: $2\rho(PP^{*})\leq\lambda(PP^{*})$. Thus, roughly speaking if
$\log\log\frac{1}{\mu_{\min}}\leq\log\frac{1}{\sqrt{\mu_{\min}}}$ (12)
then the mixing time bound based on the log-Sobolev constant is an improvement upon the mixing time bound based on the spectral gap. To make this more concrete, consider for example a Markov chain on the $d$-dimensional hypercube =$\mathcal{H}=\{-1,1\}^{d}$ with uniform stationary distribution. Then, $\mu_{\min}=2^{-d}$ and clearly Equation (12) is satisfied in this example.
Proof A.2.
of Lemma 2.2 Let the relative entropy be defined as $D(\mu^{t}:\mu):=\sum_{a\in\mathcal{A}^{N}}\mu^{t}(a)\log\frac{\mu^{t}(a)}{\mu(a)}$. Then, for a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ with irreducible transition matrix $P$ the relative entropy $D(\mu^{t}:\mu)$ decays at the following rate (Miclo, 1997):
$D(\mu^{t}:\mu)\leq(1-\rho(PP^{*}))^{t}D(\mu^{0}:\mu).$ (13)
Using Pinsker’s inequality we have that:
$\|\mu^{t}-\mu\|_{TV}\leq\sqrt{\frac{D(\mu^{t}:\mu)}{2}}\leq\sqrt{(1-\rho(PP^{*}))^{t}D(\mu^{0}:\mu)}.$ (14)
Note that $\rho(PP^{*})<1$ since $2\rho(PP^{*})\leq\lambda(PP^{*})$ by Lemma 3.1 in (Diaconis and Saloff-Coste, 1996) and for the spectral gap $\lambda(PP^{*})$ it is known that $\lambda(PP^{*})<1$ (Levin and Peres, 2017). To ensure that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon$, we derive the following lower bound on $t$:
$\sqrt{(1-\rho(PP^{*}))^{t}D(\mu^{0}:\mu)}\leq\epsilon$

### v4

@@ 3.3 Convergence time in symmetric potential games
The bound on $\rho(PP^{*})$ applies to any Markov chain whose transition matrix satisfies Equation (7). In particular, it applies to the Markov chain induced by log-linear learning since the transition matrix specified in Inequality (3) satisfies Equation (7). As this lemma is a key technical result enabling the proof of Theorem 3.1, we provide a proof in Section 4.1.
Proof (Theorem 3.1) By Cauchy-Schwarz inequality, the following holds:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\underbrace{\mathbb{E}_{a\sim\mu}[\Phi(a)]}_{\text{First term}}-2\underbrace{\|\mu^{t}-\mu\|_{TV}}_{\text{Second term}}\underbrace{\max_{a\in\mathcal{A}^{N}}\Phi(a)}_{\leq 1}.$ (8)
First term: If $\beta$ is set as in Lemma 4.1 replacing $\epsilon$ by $\epsilon/2$, then it holds that:
$\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2.$ (9)
@@
$t_{\operatorname{mix}}^{P}(\epsilon/4)\leq\frac{1}{\rho(PP^{*})}\bigg(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon}\bigg),$ (10)
where $\rho(PP^{*})$ is the log-Sobolev constant of the Markov chain induced by the transition matrix $PP^{*}$, and $P^{*}$ is the time-reversal of $P$. Therefore, if:
$t\geq\frac{1}{\rho(PP^{*})}\left(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon}\right),$ (11)
then it holds that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$. By Lemma 4.2, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
$\rho(PP^{*})\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}},$ (12)
where we used that by definition of $P$ and $\mu$ in Equations (3) and (4), respectively, $\mu_{\min}$ and $p_{\min}$ can be lower-bounded as follows:
$\mu_{\min}=\min_{a\in\mathcal{A}^{N}}\mu(a)\geq\frac{e^{-\beta}}{A^{N}}$ (13)
Plugging Inequality (12) and Inequality (13) into Inequality (11), it follows that if:
$t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\left(\log\log\frac{A^{N}}{e^{-\beta}}+2\log\frac{4}{\epsilon}\right),$
then it holds that:
$\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4.$ (14)
Combination: If $\beta$ is set as in Lemma 4.1 replacing $\epsilon$ by $\epsilon/2$ and $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\left(\log\log\frac{A^{N}}{e^{-\beta}}+2\log\frac{4}{\epsilon}\right)$, then it holds that:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\overset{(i)}{\geq}\mathbb{E}_{a\sim\mu}[\Phi(a)]-2\|\mu^{t}-\mu\|_{TV}\max_{a\in\mathcal{A}^{N}}\Phi(a)$
$\overset{(ii)}{\geq}\max_{a\in\mathcal{A}^{N}}\Phi(a)-\frac{\epsilon}{2}-\frac{2\epsilon}{4},$
$=\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
where in (i) we used Inequality (8) and (ii) follows from Inequality (9), Inequality (14), and the fact that $\Phi(\cdot)\in[0,1]$. This concludes the proof.
4.1 Proof of Lemma 4.2:
The general idea of the proof is to lower bound the log-Sobolev constant of the Markov chain $X_{t}$ with transition matrix $P$ given by Inequality (7), in terms of the log-Sobolev constant of another Markov chain, for which a lower bound is known. In particular, we make use of the following lemma.
Lemma 4.3.
[32, Corollary 2.15] Consider two Markov chains $X_{t}$ and $\hat{X}_{t}$ defined on the same state space with transition matrix $P$ and $\hat{P}$, respectively, and stationary distribution $\mu$ and $\hat{\mu}$, respectively. Then, the log-Sobolev constant $\rho(P)$ of Markov chain $X_{t}$ is lower-bounded as follows: